OpenStudy (anonymous):

answer needs to be in y=mx+b form, passes thru (1,-2) and the y intercept is 6???help

OpenStudy (anonymous):

You have two points...All you need to do is find the slope

OpenStudy (anonymous):

(0,6) and (1,-2)

OpenStudy (amistre64):

slope = (1,-2) - (0,6) = (1,-8); y/x = -8

OpenStudy (amistre64):

y = -8x +6

OpenStudy (anonymous):

Simple enough

OpenStudy (anonymous):

y=5x-7 is the answer

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hero, what is whiteboard website

OpenStudy (anonymous):

How can that be?

OpenStudy (anonymous):

AMIStre !!!! thanks a million, i found the posts

OpenStudy (anonymous):

I'll post the whiteboard site...

OpenStudy (amistre64):

yay!! :)

OpenStudy (anonymous):

idk, i got what you got too but the book says 5x-7

OpenStudy (anonymous):

you get medal , so do you hero

OpenStudy (amistre64):

5x-7 is a line that paases thru the point; but its got a yint of -7

OpenStudy (anonymous):

do you guys vaguely remember me?

OpenStudy (amistre64):

i do :)

OpenStudy (anonymous):

amistre, there is no simpler way to see all of the posts of a person, other than google site, thingy